Describe the process you would use in order to prepare 5.00 kg of an aqueous solution that is 8.00% NaCl by mass 8
1
Expert's answer
2020-04-23T13:08:29-0400

First we need to calculate the mass of pure water and NaCl:


m(NaCl) = 0.08*5.00 = 0.4 kg


m(H2O) = m(solution) - m(NaCl) = 5.00-0.4 = 4.6 kg


Next, we dissolve 0.4 kg of NaCl in 4.6 kg of water.